According to one pollster, 41% of children are afraid of the dark. Suppose that a sample of size 25 is drawn. Find the value of  , the standard deviation of the distribution of sample proportions.

Write only a number as your answer. Round to two decimal places (for example: 8.21).
Answer #1

Solution:

We are given

n = 25

p = 0.41

q = 1 – p = 1 – 0.41 = 0.59

Standard deviation of the distribution of sample proportion = Sqrt(pq/n)

Standard deviation of the distribution of sample proportion = Sqrt(0.41*0.59/25)

Standard deviation of the distribution of sample proportion = 0.098367

Answer: 0.10